How much pure water must be mixed with 6 pints of 90% developer to produce a mixture that is 17% developer?
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

How much pure water must be mixed with 6 pints of 90% developer to produce a mixture that is 17% developer?
----------------
active ingredient + active ingredient = active ingredient
0.9*6 + 0*x = 0.17(6+x)
5.4 = 1.02 + 0.17x
4.38 = 0.17x
x = 25.76 pints ( pure water must be mixed with the 90% developer)
==========================
Cheers,
Stan H.

Answer by scott8148(6628)

let x="pints of water" __ 6(90%)+x(0%)=(6+x)(17%)

multiplying by 100 to clear percentages __ 540+0=102+17x

subtracting 102 __ 438=17x __ dividing by 17 __ 438/17=x